OpenStudy (anonymous):

simplify In 9 - 4 In 3

OpenStudy (anonymous):

im stuck in the process

OpenStudy (anonymous):

start with \[4\ln(3)=\ln(3^4)\]

OpenStudy (anonymous):

then use \(\ln(a)-\ln(b)=\ln(\frac{a}{b})\) to get \[\ln(\frac{9}{81})\] then reduce

OpenStudy (anonymous):

got it thank you